# Omni2Sound: Towards Unified Video-Text-to-Audio Generation

### v1

@@ Appendix G Implementation Details
Following Stable Audio [evans2024stable], our diffusion model adopts a Diffusion Transformer (DiT) architecture within a Latent Diffusion Model (LDM) paradigm. The diffusion backbone consists of a DiT with 24 layers, 24 attention heads, and a hidden dimension of 1536. We employ cross-attention mechanisms to inject semantic conditions (e.g., FLAN-T5 and CLIP embeddings) and Adaptive Layer Normalization (AdaLN) to integrate temporal signals, as detailed in Section 4.1. Both the conditional token dimension and the global condition embedding dimension are 1024. Finally, for audio compression, we train a Variational Autoencoder (VAE) from scratch based on the wav Audio VAE architecture [evans2024stable], operating at a 16kHz sampling rate. With strides of $[4,4,4,10]$, the encoder achieves a total downsampling ratio of 640, mapping mono waveforms into a compact 64-dimensional latent space. To ensure high-fidelity reconstruction, we utilize Snake activations throughout the network.
@@
For T2A backbone pre-training, we use a large-scale corpus comprising the train set of audio datasets such as AudioCaps [kim2019audiocaps], WavCaps [mei2023wavcaps], Clotho [drossos2020clotho], AudioSet [gemmeke2017audio], VGGSound [chen2020vggsound], FSD50k [fonseca2022fsd50k], as well as music datasets including MSD [bertin2011million] and FMA [defferrard2016fma]. All audio signals are standardized to a mono-channel format at 16kHz. To accommodate fixed-size diffusion inputs, we normalize clips to a uniform 10-second duration: samples exceeding this length undergo right cropping, while shorter samples are right-padded with silence.
@@
We implement our objective evaluation metrics using the standardized AV-benchmark toolkit [hkchengrex2024avbenchmark]. All samples are generated under the same video and text conditions and evaluated in 8-second clips, following previous work [cheng2025mmaudio]. Following common practice [liu2023audioldm], we assess the quality of the generation in four critical dimensions. For Distribution Matching, we measure the similarity in feature distribution between generated and ground-truth audio. We compute the Fréchet Distance using the VGGish (FAD) [hershey2017cnn] and PaSST ($\mathrm{FD}_{\mathrm{PaSST}}$) [koutini2022passt] embeddings, as well as the Fréchet Audio Distance using PANNs (FD) [kong2020panns]. We also report the Kullback-Leibler divergence using PANNs (KL) and PaSST ($\mathrm{KL}_{\mathrm{PaSST}}$) classifiers. For Audio Quality, we assess the quality of the generation using the Inception Score [salimans2016improved], calculated with both the PANNs (IS) and PaSST ($\mathrm{IS}_{\mathrm{PaSST}}$) classifiers. For Semantic Alignment, we evaluate text-audio consistency using LAION CLAP (CLAP) [elizalde2023clap] and Microsoft CLAP (MS-CLAP) [wu2023large] scores, and video-audio alignment using ImageBind score (IB) [girdhar2023imagebind] as cosine similarity between video and audio embeddings. Finally, for Temporal Alignment, we assess audio-visual synchrony using the DS metric predicted by Synchformer [iashin2024synchformer].

### v2

@@ Appendix G Implementation Details
Following Stable Audio [13], our diffusion model adopts a Diffusion Transformer (DiT) architecture within a Latent Diffusion Model (LDM) paradigm. The diffusion backbone consists of a DiT with 24 layers, 24 attention heads, and a hidden dimension of 1536. We employ cross-attention mechanisms to inject semantic conditions (e.g., FLAN-T5 and CLIP embeddings) and Adaptive Layer Normalization (AdaLN) to integrate temporal signals, as detailed in Section 4.1. Both the conditional token dimension and the global condition embedding dimension are 1024. Finally, for audio compression, we train a Variational Autoencoder (VAE) from scratch based on the wav Audio VAE architecture [13], operating at a 16kHz sampling rate. With strides of $[4,4,4,10]$, the encoder achieves a total downsampling ratio of 640, mapping mono waveforms into a compact 64-dimensional latent space. To ensure high-fidelity reconstruction, we utilize Snake activations throughout the network.
@@
For T2A backbone pre-training, we use a large-scale corpus comprising the train set of audio datasets such as AudioCaps [24], WavCaps [36], Clotho [10], AudioSet [15], VGGSound [4], FSD50k [14], as well as music datasets including MSD [3] and FMA [9]. All audio signals are standardized to a mono-channel format at 16kHz. To accommodate fixed-size diffusion inputs, we normalize clips to a uniform 10-second duration: samples exceeding this length undergo right cropping, while shorter samples are right-padded with silence.
@@
We implement our objective evaluation metrics using the standardized AV-benchmark toolkit [7]. All samples are generated under the same video and text conditions and evaluated in 8-second clips, following previous work [7]. Following common practice [30], we assess the quality of the generation in four critical dimensions. For Distribution Matching, we measure the similarity in feature distribution between generated and ground-truth audio. We compute the Fréchet Distance using the VGGish (FAD) [19] and PaSST ($\mathrm{FD}_{\mathrm{PaSST}}$) [26] embeddings, as well as the Fréchet Audio Distance using PANNs (FD) [25]. We also report the Kullback-Leibler divergence using PANNs (KL) and PaSST ($\mathrm{KL}_{\mathrm{PaSST}}$) classifiers. For Audio Quality, we assess the quality of the generation using the Inception Score [39], calculated with both the PANNs (IS) and PaSST ($\mathrm{IS}_{\mathrm{PaSST}}$) classifiers. For Semantic Alignment, we evaluate text-audio consistency using LAION CLAP (CLAP) [11] and Microsoft CLAP (MS-CLAP) [48] scores, and video-audio alignment using ImageBind score (IB) [17] as cosine similarity between video and audio embeddings. Finally, for Temporal Alignment, we assess audio-visual synchrony using the DS metric predicted by Synchformer [22].